# Point-Level Topological Representation Learning on Point Clouds

### v2

@@ Appendix E Implementation
All experiments were run on a Apple m1 Pro chipset with 10 cores and 32 gb memory. topf and the experiments are implemented in Python and Julia. For persistent homology computations, we used gudhi (gudhi:urm) (© The gudhi developers, mit license) and Ripserer (cufar2020ripserer) (© mtsch, mit license), which is a modified Julia implementation of (Bauer2021). For the least square problems, we used the lsmr implementation of SciPy (fong2011lsmr). We used the Node2Vec python implementation https://github.com/eliorc/node2vec (© Elior Cohen, mit License) based on the Node2Vec Paper (grover2016node2vec). We used the pgeof Python package for computation of geometric features https://github.com/drprojects/point_geometric_features (© Damien Robert, Loic Landrieu, Romain Janvier, mit license). We use parts of the implementation of tpcc https://git.rwth-aachen.de/netsci/publication-2023-topological-point-cloud-clustering (© Computational Network Science Group, rwth Aachen University, mit license). We use the implementation of WSDesc Li2022WSDesc, https://github.com/craigleili/WSDesc (Lei Li, Hongbo Fu, Maks Ovsjanikov. cc by-nc 4.0 License). The idea for the implementation of the fix of the lift of the $\mathbb{Z}/3\mathbb{Z}$-representative to $\mathbb{R}$-coefficients was inspired by DREiMac, Perea2023.
@@ Appendix G How to pick the most relevant topological features
The persistent homology $P_{k}$ module in dimension $k$ is given to us as a list of pairs of birth and death times $(b_{i}^{k},d_{i}^{k})$. We can assume these pairs are ordered in non-increasing order of the durations $l_{i}^{k}=d_{i}^{k}-b_{i}^{k}$. This list is typically very long and consists to a large part of noisy homological features which vanish right after they appear. In contrast, we are interested in connected components, loops, cavities, etc. that persist over a long time, indicating that they are important for the shape of the point cloud. Distinguishing between the relevant and the irrelevant features is in general difficult and may depend on additional insights on the domain of application. In order to provide a heuristic which does not depend on any a-priori assumptions on the number of relevant features we pick the smallest quotient $q_{i}^{k}\coloneq l_{i+1}^{k}/l_{i}^{k}>0$ as the point of cut-off $\smash{N_{k}\coloneq\operatorname*{arg\,min}_{i}q_{i}^{k}}$. The only underlying assumption of this approach is that the band of ‘relevant’ features is separated from the ‘noisy’ homological features by a drop in persistence.

### v3

@@ Appendix E Implementation
All experiments were run on a Apple m1 Pro chipset with 10 cores and 32 gb memory. topf and the experiments are implemented in Python and Julia. For persistent homology computations, we used gudhi (gudhi:urm) (© The gudhi developers, mit license) and Ripserer (cufar2020ripserer) (© mtsch, mit license), which is a modified Julia implementation of (Bauer2021). For the least square problems, we used the lsmr implementation of SciPy (fong2011lsmr). We used the Node2Vec python implementation https://github.com/eliorc/node2vec (© Elior Cohen, mit License) based on the Node2Vec Paper (grover2016node2vec). We used the pgeof Python package for computation of geometric features https://github.com/drprojects/point_geometric_features (© Damien Robert, Loic Landrieu, Romain Janvier, mit license). We use parts of the implementation of tpcc https://git.rwth-aachen.de/netsci/publication-2023-topological-point-cloud-clustering (© Computational Network Science Group, rwth Aachen University, mit license). We use the implementation of WSDesc Li2022WSDesc, https://github.com/craigleili/WSDesc (Lei Li, Hongbo Fu, Maks Ovsjanikov. cc by-nc 4.0 License). We use the implementation of An Tao, https://github.com/antao97/dgcnn.pytorch of Dynamic Graph cnns wang2019dynamic, MIT license. The idea for the implementation of the fix of the lift of the $\mathbb{Z}/3\mathbb{Z}$-representative to $\mathbb{R}$-coefficients was inspired by DREiMac, Perea2023.
@@ Appendix F More Details on the Experiments
Details on experiments with DGCNN
We use the implementation of An Tao, https://github.com/antao97/dgcnn.pytorch of Dynamic Graph cnns wang2019dynamic. We use the data pretrained on the ShapeNetPart segmentation dataset, where we use the features of the second last layer. This showed the best performance. We determine the optimal hyperparameters ($k$, object class, clustering method, layer to use) for every dataset of tcbs individually and only present the results on the best parameter set. The spectral clustering always outperformed $k$-means. The authors of the original paper state that DGCNN performs very well on datasets with above $500$ points, and still is robust to downsampling beyond this point. Thus we always input the maximum available number of points to DGCNN. As described in wasserman2018topological, we scale the input point cloud to fit into a unit sphere. We pad the 2d data sets with zeroes in the third dimension. topf still outperforms DGCNN on the TCBS. We believe that this is due to the heterogeneous sampling, small sampling set size and most importantly the lack of training data for the neural network.
@@ Appendix G How to pick the most relevant topological features
The persistent homology $P_{k}$ module in dimension $k$ is given to us as a list of pairs of birth and death times $(b_{i}^{k},d_{i}^{k})$. We can assume these pairs are ordered in non-increasing order of the durations $l_{i}^{k}=d_{i}^{k}-b_{i}^{k}$. This list is typically very long and consists to a large part of noisy homological features which vanish right after they appear. In contrast, we are interested in connected components, loops, cavities, etc. that persist over a long time, indicating that they are important for the shape of the point cloud. Distinguishing between the relevant and the irrelevant features is in general difficult and may depend on additional insights on the domain of application. In order to provide a heuristic which does not depend on any a-priori assumptions on the number of relevant features we pick the smallest quotient $q_{i}^{k}\coloneq l_{i+1}^{k}/l_{i}^{k}>0$ as the point of cut-off $N_{k}\coloneq\operatorname*{arg\,min}_{i}q_{i}^{k}$. The only underlying assumption of this approach is that the band of ‘relevant’ features is separated from the ‘noisy’ homological features by a drop in persistence.